# Spectral Manifold Harmonization for Graph Imbalanced Regression

### v1

@@
Spectral Manifold Harmonization for Graph Imbalanced Regression
Brenda Nogueira Department of Computer Science and Engineering, University of Notre Dame, Notre Dame, IN, USA Gabe Gomes Department of Chemical Engineering, Carnegie Mellon University, Pittsburgh, PA, USA Meng Jiang Department of Computer Science and Engineering, University of Notre Dame, Notre Dame, IN, USA Nitesh V. Chawla Department of Computer Science and Engineering, University of Notre Dame, Notre Dame, IN, USA Nuno Moniz Department of Computer Science and Engineering, University of Notre Dame, Notre Dame, IN, USA
{onecolabstract}
@@
Graph-structured data is ubiquitous in scientific domains, where models often face imbalanced learning settings. In imbalanced regression, domain preferences focus on specific target value ranges representing the most scientifically valuable cases; we observe a significant lack of research. In this paper, we present Spectral Manifold Harmonization (SMH), a novel approach for addressing this imbalanced regression challenge on graph-structured data by generating synthetic graph samples that preserve topological properties while focusing on often underrepresented target distribution regions. Conventional methods fail in this context because they either ignore graph topology in case generation or do not target specific domain ranges, resulting in models biased toward average target values. Experimental results demonstrate the potential of SMH on chemistry and drug discovery benchmark datasets, showing consistent improvements in predictive performance for target domain ranges.
1   Introduction
@@
Graph-structured data has become increasingly important in scientific domains, particularly drug discovery, materials science, and genomics. Graph Neural Networks (GNNs) have revolutionized the modeling of such data by operating directly on graph structures, enabling more accurate predictions of molecular properties, material characteristics, and biological interactions, for example. In drug discovery alone, GNNs have demonstrated significant promise for tasks such as property prediction [18], molecular design [8], and drug-target interaction prediction [10]. The pharmaceutical industry has embraced these methods to accelerate the traditionally slow and expensive drug development pipeline, which typically costs over \$1 billion and spans more than a decade from discovery to market [14].
@@
While considerable research has targeted imbalanced classification problems in graph learning [1, 17], the regression setting has received comparatively little attention [12, 11]. Many crucial scientific problems involve predicting continuous properties where the most valuable cases are rare, e.g., in drug discovery, high-potency compounds represent a tiny fraction of the chemical space but are the most scientifically interesting [13]. Standard machine learning approaches, including GNNs, typically optimize for average performance across the entire distribution, resulting in poorly performing models on these infrequent but valuable cases. Also, existing oversampling techniques for imbalanced data fail to preserve the complex topological properties inherent in graph-structured scientific data, limiting their effectiveness in these domains.
@@
In this paper, we present Spectral Manifold Harmonization (SMH), a novel approach for tackling imbalanced regression on graph-structured data. SMH (Figure 1) operates in the graph spectral domain—the eigenspace of the graph Laplacian—to generate synthetic graph samples that preserve essential topological properties while focusing on underrepresented target distribution regions. Our approach builds on the concept of relevance in imbalanced regression [12], which maps target values to non-uniform domain preferences, assigning higher importance to specific domain ranges. SMH learns a continuous manifold of valid graph structures by establishing a mapping between target regression values and the spectral domain, allowing us to generate new samples with targeted properties.
@@
Figure 1: Visual description of the Spectral Manifold Harmonization method’s workflow.
@@ 2   Related Work
The challenge of imbalanced distributions in graph learning tasks has received increasing attention, particularly in scientific domains where rare values are critical. Recent research by [1] demonstrated that class imbalances in drug discovery datasets can be effectively addressed through techniques like oversampling and loss function manipulation when using Graph Neural Networks (GNNs). Despite these advances, most approaches operate directly in graph space rather than the spectral domain, limiting their ability to maintain global structural constraints. [4] published a comprehensive survey on spectral GNNs, highlighting their unique ability to capture global information and provide better expressiveness than spatial approaches. [15] further analyzed the theoretical expressive power of spectral GNNs, proving that they can produce arbitrary graph signals under specific conditions. However, these methods usually focus on balanced datasets, illustrating the novelty and significance of SMH.
2.1   Spectral Graph Methods
@@
Spectral graph theory has a rich history in machine learning, with applications spanning dimensionality reduction, clustering, and graph signal processing. Recent work in spectral methods includes Specformer [3], combining spectral GNNs with transformer architectures to create learnable set-to-set spectral filters, or the work by [9] to enhance the scalability of spectral GNNs without decoupling the network architecture, addressing a key limitation in previous approaches. These advanced spectral methods demonstrate improved performance on various graph learning tasks, but do not specifically target the regression setting or leverage the spectral domain for manifold harmonization in imbalanced scenarios. Our SMH method extends these ideas to regression, enabling targeted generation in underrepresented regions while maintaining global graph properties.
2.2   Manifold Learning for Structured Data
@@
Manifold learning principles underpin many approaches to generating synthetic structured data. Recently, [23] described how models can be enhanced by incorporating structured knowledge representations and latent manifold embeddings, in the context of knowledge-augmented graph machine learning for drug discovery. Similarly, [2] demonstrated that incorporating manifold information improves synthetic oversampling techniques for high-dimensional spectral data where standard approaches often fail. Our SMH approach differs from these works by explicitly modeling the regression target-to-spectrum mapping and performing manifold learning in the spectral domain, making it particularly suited for scientific applications with imbalanced regression targets.
2.3   Graph Sampling and Synthesis in Scientific Domains
@@
Due to domain-specific constraints and validity requirements, scientific applications pose unique challenges for graph-based methods. [20] provided a comprehensive bibliometric analysis of GNN applications in drug discovery, showing significant growth in this area and highlighting the need for methods to handle the inherent data imbalances in these domains. Similarly, [6] addressed the challenge of overconfident errors in molecular property classification, demonstrating the importance of uncertainty quantification in imbalanced datasets. These approaches focus primarily on classification rather than regression tasks, and do not specifically utilize spectral representations to address imbalance.
@@
On regression tasks, a review on GNNs for predicting synergistic drug combinations [22] noted that graph-based models often suffer from imbalanced data distributions, affecting their performance. They emphasized the need for methods to handle such imbalances to improve predictive accuracy effectively. Our SMH method offers a domain-agnostic approach that incorporates scientific validity constraints while focusing on generating underrepresented regions of the target distribution, bridging critical gaps in existing methodologies for imbalanced regression on graph-structured data in scientific applications.
3   Methods: Spectral Manifold Harmonization
@@
Our Spectral Manifold Harmonization (SMH) method addresses imbalanced regression on graph-structured data by learning to generate synthetic graph samples in underrepresented regions of the target distribution while preserving their topological properties. The key insight is that operating in the graph spectral domain allows us to construct a continuous manifold of valid graph structures, making it possible to sample new graphs with targeted properties. SMH integrates the concept of relevance from recent work on imbalanced regression [12, 13] and consists of five main components (Figure 1): we first transform graphs into their spectral representation, learn how target values map to this spectral space with emphasis on relevant regions, model the manifold of valid spectral representations, strategically sample from underrepresented areas, and finally transform back to generate new graph instances that address the imbalance problem.
3.1   Graph Spectral Representation and Relevance Concept
@@
A key concept in addressing imbalanced regression is relevance, which maps target values to non-uniform domain preferences [12]. In this context, a continuous, domain-dependent relevance function $\phi(Y):\mathcal{Y}\rightarrow[0,1]$ expresses the application-specific bias concerning the target variable $\mathcal{Y}$. A domain expert ideally defines the relevance function for the specific task where the expert inputs information on the available target value-relevance pairs, i.e., which value is considered low or high-relevance. When this information is unavailable, the function can be interpolated from boxplot-based statistics where extreme values are considered high-relevance and the distribution median is considered the lowest point of relevance.
3.2   Relevance-Guided Target-to-Spectrum Mapping
@@
$f_{\theta}(y)=\mathbf{W}_{L}\cdot\sigma(\mathbf{W}_{L-1}\cdot\sigma(\cdots\sigma(\mathbf{W}_{1}\cdot y+\mathbf{b}_{1})\cdots)+\mathbf{b}_{L-1})+\mathbf{b}_{L}$ (1)
@@
$\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\phi(y_{i})\cdot\|\mathbf{s}_{i}-f_{\theta}(y_{i})\|^{2}+\alpha\cdot\Omega(\theta)$ (2)
@@
3.3   Manifold Learning in Spectral Space
@@
$p(\mathbf{s}|y)=\mathcal{N}(\mu(y),\Sigma(y))$ (3)
@@
$\Sigma(y)=\sum_{i=1}^{N}w_{i}(y)\cdot(\mathbf{s}_{i}-\mu(y))(\mathbf{s}_{i}-\mu(y))^{T}$ (4)
@@
3.4   Constrained Sampling for Underrepresented Regions
@@
3.5   Inverse Spectral Transformation
@@ 5. RQ5
4.1   Methods
We converted the SMILES into a networkx [7] graph to build the spectral manifold harmonization space. Then, we used XGBoost to train a model to predict the eigenvalues from a given target. For property prediction, we then convert the networkx graph format for PyTorch Geometric data format and input it in a Graph Isomorphism Network (GIN) [19], as it has emerged as a powerful tool for graph-based machine learning tasks due to of its capability to effectively differentiate between different graph structures, using MSE as a loss function. The hyperparameter is presented in Appendix A with a 5-fold cross-validation. We also compared our relevance-guided target-to-spectrum transformation and constrained sampling approach with SMOGN [5]. For the Spectral+SMOGN baseline, we first compute the spectral representation as described in Section 3.1, and then apply the SMOGN method. The inverse transformation used for decoding remains the same. The code is available at: https://github.com/brendacnogueira/smh-graph-imbalance.git. To compare with a pre-trained model, we used HiMol [21], which is a framework to learn molecule graph representation for property prediction.
4.2   Data
Our experimental evaluation focuses on molecular data, using regression tasks from MoleculeNet [16]: ESOL, FreeSolv, and Lipophilicity (Lipo). The datasets are briefly described in Table 1. The datasets exhibit a long-tailed distribution toward the lower end of the property range, and we define our relevance function to assign higher importance to these.
Table 1: Summary of Molecular Property Datasets
@@ 5   Results and Discussion
5.1   Synthetic Generated Graphs
@@
Figure 2: Illustration of graphs selected for augmentation and the corresponding synthetic graphs generated using the approach described in Section 3, for the ESOL dataset.
A comparison of the mean and standard deviation of node and edge counts between the original and synthetic graphs is provided in Table 2. The number of nodes remains nearly identical across SHM and SMOGN. Minor differences are observed in the number of edges and graph density, with the synthetic graphs showing slightly lower mean values. However, these differences remain within an acceptable range, supporting the validity of the generated graphs (RQ1). Further validation on the generated graphs can be addressed.
Table 2: Comparison of Mean and Standard Deviation of Node and Edge Counts Between Original and Synthetic Graphs.
Data set Node Edge Density
FreeSolv Baseline $8.7\pm 4.19$ $8.4\pm 4.79$ $0.3\pm 0.15$
SMH $10.1\pm 3.10$ $5.0\pm 4.63$ $0.1\pm 0.12$
Spectral+SMOGN $10.6\pm 2.72$ $10.8\pm 3.69$ $0.2\pm 0.07$
ESOL Baseline $13.3\pm 6.93$ $13.7\pm 8.00$ $0.2\pm 0.13$
SMH $20.5\pm 4.40$ $11.3\pm 11.42$ $0.1\pm 0.06$
SMOGN $20.8\pm 3.43$ $22.6\pm 3.71$ $0.1\pm 0.02$
Lipo Baseline $27.1\pm 7.34$ $29.5\pm 8.12$ $0.1\pm 0.03$
SMH $23.8\pm 12.69$ $14.5\pm 14.28$ $0.1\pm 0.05$
Spectral+SMOGN $23.7\pm 16.59$ $30.3\pm 23.64$ $0.1\pm 0.08$
5.2   Model Performance
@@
Table 3: Experimental results for the FreeSolv, ESOL, and LIPO datasets, using the SERA, MAE, RMSE, and $R^{2}$ evaluation metrics. Arrows signal the direction for best results, also noted in bold.
@@
Figure 3: Distribution of train dataset with and without synthetic augmentation, along with the improvements for each part of the test set domain, for each dataset.
6   Conclusion

### v2

@@
Abstract.
@@
Graph-structured data is ubiquitous in scientific domains, where models often face imbalanced learning settings. In imbalanced regression, domain preferences focus on specific target value ranges that represent the most scientifically valuable cases; however, we observe a significant lack of research regarding this challenge. In this paper, we present Spectral Manifold Harmonization (SMH), a novel approach to address imbalanced regression challenges on graph-structured data by generating synthetic graph samples that preserve topological properties while focusing on the most relevant target distribution regions. Conventional methods fail in this context because they either ignore graph topology in case generation or do not target specific domain ranges, resulting in models biased toward average target values. Experimental results demonstrate the potential of SMH on chemistry and drug discovery benchmark datasets, showing consistent improvements in predictive performance for target domain ranges. Code is available at https://github.com/brendacnogueira/smh-graph-imbalance.git.
Graph Neural Networks, Spectral Graph Theory, Imbalanced Regression, Scientific Machine Learning, Drug Discovery, Materials Science
copyright: acmlicensedjournalyear: 2025conference: Workshop on Machine Learning on Graphs in the Era of Generative Artificial Intelligence; August 6, 2025; Toronto, ON, Canadaisbn: ccs: Computing methodologies Spectral methodsccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Supervised learning by regression
1. Introduction
@@
Graph-structured data has become increasingly important in scientific domains, particularly drug discovery, materials science, and genomics. Graph Neural Networks (GNNs) have revolutionized the modeling of such data by operating directly on graph structures, enabling more accurate predictions of molecular properties, material characteristics, and biological interactions, for example. In drug discovery alone, GNNs have demonstrated significant promise for tasks such as property prediction (Xiong et al., 2020), molecular design (Jin et al., 2018), and drug-target interaction prediction (Lim et al., 2019). The pharmaceutical industry has embraced these methods to accelerate the traditional drug development pipeline, which typically costs over \$1 billion and spans more than a decade from discovery to market (Vamathevan et al., 2019).
@@
While considerable research has targeted imbalanced classification problems in graph learning (Almeida et al., 2024; Xia et al., 2024), the regression setting has received little attention (Ribeiro and Moniz, 2020; Liu et al., 2023). Crucial scientific problems involve predicting continuous properties where the most valuable cases are rare, e.g., in drug discovery, high-potency compounds represent a tiny fraction of the chemical space but are the most scientifically interesting (Silva et al., 2022). Standard machine learning approaches, including GNNs, typically optimize for average performance across the entire distribution, resulting in models that perform poorly on these traditionally infrequent but valuable cases. Additionally, existing oversampling techniques for imbalanced data often fail to preserve the complex topological properties inherent in graph-structured scientific data, thereby limiting their effectiveness in these domains.
@@
In this paper, we present Spectral Manifold Harmonization (SMH), a novel approach for tackling imbalanced regression (Ribeiro and Moniz, 2020; Moniz et al., 2017; Ma et al., 2024) on graph-structured data. SMH (Figure 1) operates in the graph spectral domain—the eigenspace of the graph Laplacian—to generate synthetic graph samples that preserve essential topological properties while focusing on underrepresented target distribution regions. Building on the concept of relevance in imbalanced regression (Ribeiro and Moniz, 2020), which maps target values to non-uniform domain preferences, SMH learns a continuous manifold of valid graph structures by relating target values and the spectral domain, allowing the generation of new samples with targeted properties.
@@
Figure 1. Visual description of the Spectral Manifold Harmonization method’s workflow.
[…]
@@ 2. Related Work
The challenge of imbalanced distributions in graph learning tasks has received increasing attention, particularly in scientific domains where rare values are critical. Recent research by Almeida et al. (2024) demonstrates that imbalanced learning in drug discovery datasets can be tackled with techniques such as oversampling and loss function manipulation when using Graph Neural Networks (GNNs). Despite these advances, most approaches operate directly in graph space rather than the spectral domain, limiting their ability to maintain global structural constraints. Bo et al. (2023b) published a comprehensive survey on spectral GNNs, highlighting their unique ability to capture global information and provide better expressiveness than spatial approaches. Wang and Zhang (2022) further analyzed the theoretical expressive power of spectral GNNs, proving that they can produce arbitrary graph signals under specific conditions. However, these methods focus on balanced and classification datasets, illustrating the novelty and significance of SMH.
2.1. Spectral Graph Methods
@@
Spectral graph theory has a rich history in machine learning, with applications spanning dimensionality reduction, clustering, and graph signal processing. Recent work in spectral methods includes Specformer (Bo et al., 2023a), combining spectral GNNs with transformer architectures to create learnable set-to-set spectral filters, or the work by (Li et al., 2025) to enhance the scalability of spectral GNNs without decoupling the network architecture, addressing a key limitation in previous approaches. These advanced spectral methods demonstrate improved performance on various graph learning tasks, but do not specifically target the regression setting or leverage the spectral domain for manifold harmonization in imbalanced scenarios. Our SMH method extends these ideas to regression, enabling targeted generation in underrepresented regions while maintaining global graph properties.
2.2. Manifold Learning for Structured Data
@@
Manifold learning principles underpin many approaches to generating synthetic structured data. Recently, Zhong et al. (2024) described how models can be enhanced by incorporating structured knowledge representations and latent manifold embeddings, in the context of knowledge-augmented graph machine learning for drug discovery. Similarly, Baumgartner et al. (2023) demonstrated that incorporating manifold information improves synthetic oversampling techniques for high-dimensional spectral data where standard approaches often fail. Our SMH approach differs from these works by explicitly modeling the regression target-to-spectrum mapping and performing manifold learning in the spectral domain, making it particularly suited for scientific applications with imbalanced regression targets.
2.3. Graph Sampling and Synthesis in Scientific Domains
@@
Due to domain-specific constraints and validity requirements, scientific applications pose unique challenges for graph-based methods. Yao et al. (2024) provided a comprehensive bibliometric analysis of GNN applications in drug discovery, showing significant growth in this area and highlighting the need for methods to handle the inherent data imbalances in these domains. Similarly, Fan et al. (2024) addressed the challenge of overconfident errors in molecular property classification, demonstrating the importance of uncertainty quantification in imbalanced datasets. These approaches focus primarily on classification rather than regression tasks, and do not specifically utilize spectral representations to address imbalance.
@@
On regression tasks, a review on GNNs for predicting synergistic drug combinations (Zhang and Tu, 2023) noted that graph-based models often suffer from imbalanced data distributions, affecting their performance. They emphasized the need for methods to handle such imbalances to improve predictive accuracy effectively. Our SMH method offers a domain-agnostic approach that incorporates scientific validity constraints while focusing on generating underrepresented regions of the target distribution, bridging critical gaps in existing methodologies for imbalanced regression on graph-structured data in scientific applications.
3. Methods: Spectral Manifold Harmonization
@@
Our Spectral Manifold Harmonization (SMH) method addresses imbalanced regression on graph-structured data by learning to generate synthetic graph samples in underrepresented regions of the target distribution while preserving their topological properties. The key insight is that operating in the graph spectral domain allows us to construct a continuous manifold of valid graph structures, making it possible to sample new graphs with targeted properties. SMH integrates the concept of relevance from recent work on imbalanced regression (Ribeiro and Moniz, 2020; Silva et al., 2022) and consists of five main components (Figure 1): we first transform graphs into their spectral representation, learn how target values map to this spectral space with emphasis on relevant regions, model the manifold of valid spectral representations, strategically sample from underrepresented areas, and finally transform back to generate new graph instances that address the imbalance problem.
3.1. Graph Spectral Representation and Relevance Concept
@@
A key concept in addressing imbalanced regression is relevance, which maps target values to non-uniform domain preferences (Ribeiro and Moniz, 2020). In this context, a continuous, domain-dependent relevance function $\phi(Y):\mathcal{Y}\rightarrow[0,1]$ expresses the application-specific bias concerning the target variable $\mathcal{Y}$. A domain expert ideally defines the relevance function for the specific task where the expert inputs information on the available target value-relevance pairs, i.e., which value is considered low or high-relevance. When this information is unavailable, the function can be interpolated from boxplot-based statistics where extreme values are considered high-relevance and the distribution median is considered the lowest point of relevance.
3.2. Relevance-Guided Target-to-Spectrum Mapping
@@
(1) $f_{\theta}(y)=\mathbf{W}_{L}\cdot\sigma(\mathbf{W}_{L-1}\cdot\sigma(\cdots\sigma(\mathbf{W}_{1}\cdot y+\mathbf{b}_{1})\cdots)+\mathbf{b}_{L-1})+\mathbf{b}_{L}$
@@
(2) $\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\phi(y_{i})\cdot\|\mathbf{s}_{i}-f_{\theta}(y_{i})\|^{2}+\alpha\cdot\Omega(\theta)$
@@
3.3. Manifold Learning in Spectral Space
@@
(3) $p(\mathbf{s}|y)=\mathcal{N}(\mu(y),\Sigma(y))$
@@
(4) $\Sigma(y)=\sum_{i=1}^{N}w_{i}(y)\cdot(\mathbf{s}_{i}-\mu(y))(\mathbf{s}_{i}-\mu(y))^{T}$
@@
3.4. Constrained Sampling for Underrepresented Regions
@@
3.5. Inverse Spectral Transformation
@@ 4. Experiments
4.1. Methods
We converted the SMILES into a networkx (Hagberg and Conway, 2020) graph to build the spectral manifold harmonization space. Then, we used XGBoost to train a model to predict the eigenvalues from a given target. For property prediction, we then convert the networkx graph format for PyTorch Geometric data format and input it in a Graph Isomorphism Network (GIN) (Xu et al., 2018), a powerful tool for graph-based machine learning tasks due to of its capability to effectively differentiate between different graph structures, using MSE as a loss function. The hyperparameter is presented in Appendix A with a 5-fold cross-validation. We also compared our relevance-guided target-to-spectrum transformation and constrained sampling approach with SMOGN (Branco et al., 2017). For the Spectral+SMOGN baseline, we first compute the spectral representation as described in Section 3.1, and then apply the SMOGN method. The inverse transformation used for decoding remains the same. To compare with a pre-trained model, we used HiMol (Zang et al., 2023), which is a framework to learn molecule graph representation for property prediction.
4.2. Data
Our experimental evaluation focuses on molecular data, using regression tasks from MoleculeNet (Wu et al., 2018): ESOL, FreeSolv, and Lipophilicity (Lipo). The datasets are briefly described in Table 1. The datasets exhibit a long-tailed distribution toward the lower end of the property range, and we define our relevance function to assign higher importance to these.
Table 1. Summary of Molecular Property Datasets
@@ 5. Results and Discussion
5.1. Synthetic Generated Graphs
@@
Figure 2. Illustration of graphs selected for augmentation and the corresponding synthetic graphs generated using the approach described in Section 3, for the ESOL dataset.
A comparison of the mean and standard deviation of node and edge counts between the original and synthetic graphs is provided in Figure 3. The number of nodes remains very similar across sets. Minor differences are observed in the number of edges and graph density, with the synthetic graphs showing slightly higher mean values. However, these differences remain within an acceptable range, supporting the validity of the generated graphs (RQ1). Further validation on the generated graphs can be addressed.
Figure 3. Comparison of Mean and Standard Deviation of Node and Edge Counts Between Original and Synthetic Graphs.
5.2. Model Performance
@@
Table 2. Experimental results for the FreeSolv, ESOL, and LIPO datasets, using the SERA, MAE, RMSE, and $R^{2}$ evaluation metrics. Arrows signal the direction for best results, also noted in bold.
@@
Figure 4. Distribution of train dataset with and without synthetic augmentation, along with the improvements for each part of the test set domain, for each dataset.
6. Conclusion